Since the book bag slide on the car seat, what is
Answer #1

V0x = 8.66 m/s


V0y = 5 m/s


at the maximum height

vertical component vy = 0


displacement y = H

vy^2 - v0y^2 = 2*ay*g


v0y^2 = 2*g*H


H = v0y^2/(2*g)

H = 5^2/(2*9.8)

H = 1.27 m


==================

after falling on the ground


along vertical


total displacement y = 0

y = v0y*T + (1/2)*ay*T^2

T = 2*V0y/g


time T = 2*v0y/g = 2*5/9.8 = 1.02 s

==========


along horizantal

ax = 0

X = V0x*T + (1/2)*ax*T^2


X = V0x*T

X = V0x*2*V0y/g


X = 8.66*2*5/9.8 = 8.84 m


----------